# gLSTM: Mitigating Over-Squashing by Increasing Storage Capacity

### v1

@@ 1 Introduction
Over-squashing was initially identified by Alon & Yahav (2021) as a problem of capacity; the compression of information from a node’s receptive field into a fixed-size vector. This was linked with depth and long-range dependencies, since on many graphs receptive fields grow exponentially with depth. Later work (Topping et al., 2022; Di Giovanni et al., 2023a) identified that this bottleneck could also result in low sensitivity as measured by the node Jacobian, linking graph topology and aspects of model architecture via an upper bound on this Jacobian. This low sensitivity arises due to repeated degree normalization and application of a contractive nonlinearity over many layers. We argue (Sections 3 and 5.2) that these two issues – limited information storage capacity, and low sensitivity – are distinct. This work revives the capacity viewpoint of over-squashing which has received little attention since the work of Alon & Yahav (2021), and studies it in isolation from sensitivity issues. We believe this focus on capacity not only provides a more complete understanding of over-squashing but also highlights new directions to mitigate it.
To combat over-squashing, existing research has focused on ameliorating topological bottlenecks through rewiring (Gasteiger et al., 2019; Gutteridge et al., 2023; Nguyen et al., 2023) and controlling the flow of information (Bresson & Laurent, 2017; Finkelshtein et al., 2024; Errica et al., 2025) – both designed to reduce bottlenecks in the computational graph. However, this bottleneck is only an issue if it (1) results in low sensitivity or (2) saturates the storage capacity. We target the latter failure mode: adapting the MPNN architecture to improve its ability to store and retrieve information. Framing over-squashing as a capacity limitation that can be addressed at the architecture level exposes a previously unexplored path, and our results validate this direction.
@@
Our main contributions are as follows. In Section 3, we re-characterize over-squashing as a combination of two distinct issues: saturating capacity and low sensitivity, which we term capacity over-squashing and sensitivity over-squashing respectively. We discuss in Section 3.1 the pitfalls of widely used over-squashing tasks, which either fail to evaluate capacity at all, or evaluate the two issues in tandem and are thus unable to separate their effects. In Section 3.2, we introduce a novel synthetic task, which to our knowledge is the first that measures capacity over-squashing in isolation. In Section 4, we present a new MPNN architecture based on the recent xLSTM architecture (Beck et al., 2024), which uses associative memory to increase capacity, explicitly targeting this capacity over-squashing viewpoint. Section 5 demonstrates that this architecture performs well on our synthetic capacity task and a range of real-world benchmarks, and Section 5.2 demonstrates empirically that capacity over-squashing can occur separately from sensitivity over-squashing.
@@ 2 Background and Related Work
GNNs are functions $f_{\bm{\theta}}\!:\!(\mathcal{G},\{{\bm{x}}_{v}\})\mapsto{\bm{y}}$ with parameters $\bm{\theta}$, trained via gradient descent to predict node- or graph-level labels $\bm{y}$}. These models typically take the form of MPNNs, which compute latent representations by composing $L$ layers of the following “message passing” node-wise operation:
@@
Recent work Beck et al. (2024; 2025) introduced xLSTM, a development of the original LSTM Hochreiter & Schmidhuber (1997) architecture that resulted in a performant recurrent neural network capable of language modeling. Of relevance to our work are the following LSTM limitations that xLSTM aims to address: the inability to “revise storage decisions” and the limited storage capacity of the scalar cell states. The first of these issues is addressed through modifying the original LSTM gating to use exponential activation functions rather than sigmoid functions. The second is addressed by introducing associative memory, updated using an outer product update rule equivalent to that of FWPs to store keys and values (see Appendix A for more details).
3 The Two Faces of Over-Squashing
@@
We argue that there are two distinct problems arising from bottlenecks in MPNNs: reduced sensitivity (sensitivity over-squashing) and saturating storage capacity (capacity over-squashing). While due to the influential paper of Topping et al. (2022) the sensitivity viewpoint on over-squashing has thus far been the predominant approach in the literature; in this work, we seek to revisit the storage capacity viewpoint and investigate how this issue can be avoided. We define storage capacity as the amount of information that can be stored in a node’s representation for later use: a representation is saturated when it is unable to store any more information.
@@ 3.1 Existing Over-Squashing Tasks do not (Only) test Capacity
Figure 1: Computational graphs. Left: RingTransfer (Di Giovanni et al., 2023a). Middle: Tree-NeighborsMatch (Alon & Yahav, 2021). Right: NAR, introduced in Section 3.2. Nodes with informative features are green, background gray. Output of the red node is trained to solve the task.
@@
An instructive way of contrasting sensitivity against capacity is via synthetic tasks. The most common synthetic tasks currently used to assess over-squashing are the RingTransfer tasks of Di Giovanni et al. (2023a). The goal of these tests is for a MPNN to ‘transfer’ features contained at a target node to a source node, across a large graph distance. Various graphs are tested, in particular a ring of nodes, but the common feature is that there exists a long shortest-path from the source to target node, so the information must be passed through many nodes. All of these exhibit an exponentially growing receptive field of at least $2^{k}$ for $k$ layers, since each node is connected to at least two others; repeated aggregation over neighborhoods and application of MPNN layers and nonlinearities makes this a good test of the sensitivity-based view of over-squashing.
However, this task is particularly ill-equipped to test the issue of storage capacity, as the only relevant information in the graph is that of the target node, and all intermediate nodes are assigned constant vectors of ones. In this way, there is only a single node’s representation worth of information to be transferred. It is unclear how much this task measures behavior found in real-world tasks, as over-squashing will not comprise sensitivity issues in isolation: exponentially growing receptive fields will not be padded by nodes with identical representations containing no new information. Figure 2 (left) visualizes the computational graph of RingTransfer, demonstrating that it is dominated by nodes containing no information.
Alon & Yahav (2021) introduced the Tree-NeighborsMatch task to measure capacity via propagating information from the leaf nodes of a variable-size binary tree. It shares similarities with the task we introduce in Section 3.2 in that it controls the amount of information that is forced into a bottleneck of a single node representation. However, it propagates this information through a deep binary tree, requiring variable-depth MPNNs to solve this task. This depth significantly worsens sensitivity: we visualize Jacobian norms of a GCN acting on a deep binary tree vs a single layer tree with matching leaf counts in Figure 2, demonstrating that this sensitivity drops off far faster for deep GCNs. This is unsurprising given the bound of Equation 3, showing that the deep GCN must additionally contend with “model” squashing terms of nonlinearity and weight contraction that scale exponentially with depth. Therefore we expect performance degradation trends to be due to both 1) saturating capacity and 2) low sensitivity; deep tasks such as Tree-NeighborsMatch are impacted by both over-squashing issues, rather than isolating the issue of capacity.
@@ 3.2 Neighbor Associative Recall: Isolating Storage Capacity
For a fixed neighborhood size $N$ we define a fixed set of keys and values, $N=|{\mathbb{K}}|=|{\mathbb{V}}|$, and a pair of learned vector embedding functions ${\bm{e}}_{k}:{\mathbb{K}}\to{\mathbb{R}}^{d_{\text{emb}}}$, ${\bm{e}}_{v}:{\mathbb{V}}\to{\mathbb{R}}^{d_{\text{emb}}}$ for embedding dimension $d_{\text{emb}}$. Each of the neighbor nodes $n$ has a different assigned key $k_{n}\in{\mathbb{K}}$, and also a value $v_{n}\in{\mathbb{V}}$, randomly sampled with replacement. The input feature vector of these nodes is a concatenation of the two learned embeddings ${\bm{x}}_{n}=[{\bm{e}}_{k}(k_{n});{\bm{e}}_{v}(v_{n})]\in{\mathbb{R}}^{2d_{\text{emb}}}$. The intermediate node and central node both have zero-valued feature vectors. Associated with the query node $q$ is a randomly sampled key-value node $m$; the input feature vector for the query node consists of the corresponding key embedding concatenated with a vector of zeroes, ${\bm{x}}_{q}=[{\bm{e}}_{k}(k_{m});\bm{0}]\in{\mathbb{R}}^{2d_{\text{emb}}}$. The model is trained such that the central node must predict the value $v_{m}$ associated with the sampled key node. Training is via cross-entropy loss where the target of the central node is a one-hot vector corresponding to a fixed ordering of $\mathbb{V}$}. This approach can be viewed as a graph adaptation of the sequence associative recall task of Schlag et al. (2021). Results are presented in Section 5.1. An alternative formulation of this task with a regression target is discussed in Section B.6.
@@
Prior work on over-squashing has focused almost exclusively on mitigating sensitivity issues, often through graph rewiring (Gasteiger et al., 2019; Gutteridge et al., 2023; Nguyen et al., 2023). By contrast, we find no studies that tackle capacity over-squashing directly. Motivated by memory-capacity gains in sequence models (Ba et al., 2016; Beck et al., 2024), we introduce associative memory into an MPNN architecture to explicitly enlarge its information-storage capacity. We can specifically measure this increased capacity in the graph setting using the NAR task, introduced above. We further introduce the gating scheme of Beck et al. (2024) to investigate its efficacy in the graph setting, given strong sequence modeling performance. Since these adaptations are inspired in part by their successful use in xLSTM, we refer to our related graph architecture as gLSTM.
@@
The modified gLSTM update equations are given below. Highlighted in blue are the differences to xLSTM. Biases correspond exactly to xLSTM (Appendix A) and are omitted for clarity.
@@
Figure 4: gLSTM block structure. Gate layers shown in orange, query/key/value computation in dark blue. Aggr. is short for aggregation, and represents (K-hop) aggregation across neighborhoods. Symbols $\odot,\otimes,+,\cdot$ denote respectively Hadamard product, outer product, vector addition, matrix multiplication.
@@
This aggregation scheme appears to greatly improve gLSTM performance: our synthetic task in Section 5.1 significantly benefits from this aggregation scheme, and the ablations in Section B.2 demonstrate that it improves performance in all but one of the tested benchmarks. We hypothesize that – in addition to providing a highly connected computational graph that lessens over-squashing sensitivity bottleneck issues – this is because it also provides an extremely useful inductive bias for the recall mechanism of gLSTM. Information that has previously been stored in the associative memory is not then included in later message passing rounds, and later nodes are able to query this memory in isolation.
@@ 5.1 Neighbor Associative Recall
We train various models on NAR with varying neighbor count $N$, with results shown in Figure 5(a). Throughout this section we compare gLSTM using K-hop aggregation to GCN using standard aggregation, since gLSTM performs significantly better in this task when using K-hop aggregation whereas GCN performance is harmed by K-hop. We present additional results in Section B.5 where we separate by aggregation method and include results for a larger number of models. A comparison of the number of trainable parameters is shown in Figure 5(b). Fair comparison between matrix and vector memory is nontrivial, so we select these parameter counts to “favor” GCN.
@@
Figure 5: Test-set mean Accuracy (standard deviation shaded) for the NAR task, for gLSTM and GCN models with various hidden dimensions shown in Figure 5(a), number of trainable parameters in Figure 5(b). Note that gLSTM uses K-hop aggregation here, whereas GCN does not; see Section B.5 for separated performance by aggregation strategy.
@@ 5.2 How does Capacity Relate to Sensitivity?
We see therefore that sensitivity, as measured by the Jacobian norm, does not correlate with NAR performance. Given that NAR performance degradation is due to capacity over-squashing, we therefore observe that capacity over-squashing can occur without sensitivity over-squashing. This is clear from the fact that 1) sensitivity increases consistently for GCN models above $N=16$ to the point where it matches initial sensitivity, despite no increase in performance and 2) sensitivity for gLSTM tends to carry on increasing beyond where performance starts to degrade. We note these trends – as with all observations we make in this section – hold true for the NAR regression task in Section B.6.1.
@@ 5.3 Long Range Benchmarks
Table 1: Mean and standard deviation of $\log_{10}(\text{MSE})$, averaged over 4 random weight initializations on the GPP tasks from Gravina et al. (2023), from which we report baselines. See Section B.1 for discussion of baseline choice. Top score in bold, second underlined. Lower is better.
@@
gLSTM achieves comfortably state of the art results on the Diameter and Eccentricity GPP tasks, and very strong performance on SSSP; notably SSSP is the only tested task in which k-hop decreases performance. LRGB results show that gLSTM achieves strong performance in Peptides-Func but relatively weak performance on Peptides-Struct. We hypothesize that the weaker performance on Peptides-Struct may be due to long-range interactions being less relevant for this task, which is very effectively solved by a few-layer GCN. See Section B.2 for gLSTM ablations on these benchmarks and Section B.3 for details around hyperparameters used.

### v2

@@ 1 Introduction
Over-squashing was initially identified by Alon & Yahav (2021) as a problem of compressing information from a node’s receptive field into a fixed-size vector. This was linked with depth and long-range dependencies, since receptive fields tend to grow exponentially with depth. Later work (Topping et al., 2022; Di Giovanni et al., 2023a) identified that this bottleneck could also result in low sensitivity as measured by the node Jacobian, linking graph topology and aspects of model architecture via an upper bound on this Jacobian. This low sensitivity arises due to repeated degree normalization and application of a contractive nonlinearity over many layers. Arnaiz-Rodriguez & Errica (2025) suggest that these two descriptions of over-squashing are not the same, and that Alon & Yahav (2021) define it as a problem of computational graph bottlenecks, while later work often defines it as a problem of topological bottlenecks. We discuss this separation and its relation to our work in Appendix A.
Instead of separating by issues of computational tree structure and bottlenecks, we suggest an alternative separation by resultant failure mode: limited information storage capacity, and low sensitivity. In light of this, we highlight another divergence in the literature: the work of Alon & Yahav (2021) implicitly described over-squashing as a capacity problem, and later work re-framed it as a problem of sensitivity. Focusing on the failure modes themselves not only allows us to revisit the issue of capacity originally discussed in Alon & Yahav (2021), but also to more directly motivate benchmarks and remedies, both of which we discuss in this paper. We further contextualize this issue of capacity by studying it in isolation from sensitivity issues. We believe this focus not only provides a more complete understanding of over-squashing but also highlights new directions to mitigate it.
To combat over-squashing, existing research has focused on ameliorating topological bottlenecks through rewiring (Gasteiger et al., 2019; Gutteridge et al., 2023; Nguyen et al., 2023) and controlling the flow of information (Bresson & Laurent, 2017; Finkelshtein et al., 2024; Errica et al., 2025) – targeting topological and general computational bottlenecks respectively. However, these bottlenecks are only an issue if they harm performance in some way: in this work we discuss issues of (1) reduced sensitivity and (2) saturated storage capacity. Our proposed architecture in Section 4 targets the latter failure mode: adapting the MPNN architecture to improve its ability to store and retrieve information. Framing over-squashing as a capacity limitation that can be addressed at the architecture level exposes a previously unexplored path, and our results validate this direction.
@@
Our main contributions are as follows. In Section 3, we re-characterize over-squashing into two distinct failure modes: saturating capacity and low sensitivity, which we term capacity over-squashing and sensitivity over-squashing respectively. We discuss in Section 3.1 the pitfalls of widely used over-squashing tasks, which either fail to evaluate capacity at all, or evaluate the two issues in tandem and are thus unable to separate their effects. In Section 3.2, we introduce a novel synthetic task, which to our knowledge is the first that measures capacity over-squashing in isolation. In Section 4, we present a new MPNN architecture based on the recent xLSTM architecture (Beck et al., 2024), which uses associative memory to increase capacity, explicitly targeting this capacity over-squashing viewpoint. Section 5 demonstrates that this architecture performs well on our synthetic capacity task and a range of real-world benchmarks, and Section 5.2 demonstrates empirically that capacity over-squashing can occur separately from sensitivity over-squashing.
@@ 2 Background and Related Work
GNNs are functions $f_{\bm{\theta}}\!:\!(\mathcal{G},\{{\bm{x}}_{v}\})\mapsto{\bm{y}}$ with parameters $\bm{\theta}$, trained via gradient descent to predict node- or graph-level labels $\bm{y}$}. These models typically take the form of MPNNs, which compute latent representations by composing $L$ layers of the following node-wise operation:
@@
Recent work Beck et al. (2024; 2025) introduced xLSTM, a development of the original LSTM Hochreiter & Schmidhuber (1997) architecture that resulted in a performant recurrent neural network capable of language modeling. Of relevance to our work are the following limitations that xLSTM aims to address: the inability to “revise storage decisions” and the limited storage capacity of the scalar cell states. The first of these is addressed through modifying the original LSTM gating to use exponential activation functions. The second is addressed by introducing associative memory, updated using an outer product update rule equivalent to that of FWPs to store keys and values (see Appendix B for more details).
3 The Two Failure Modes of Over-Squashing
@@
We argue that there are two distinct problems arising from bottlenecks in MPNNs: reduced sensitivity (sensitivity over-squashing) and saturating storage capacity (capacity over-squashing). Due to the influential paper of Topping et al. (2022) the sensitivity viewpoint on over-squashing has thus far been the predominant approach in the literature; in this work, we seek to revisit the storage capacity viewpoint and investigate how this issue can be avoided. We define storage capacity as the amount of information that can be stored in a node’s representation for later use: a representation is saturated when it is unable to store any more information.
@@ 3.1 Existing Over-Squashing Tasks do not (Only) test Capacity
Figure 1: Computational graphs. Left: RingTransfer (Di Giovanni et al., 2023a). Middle: Tree-NeighborsMatch (Alon & Yahav, 2021). Right: NAR, introduced in Section 3.2. Nodes with informative features are green, background gray. Red node is trained to solve the task.
@@
An instructive way of contrasting sensitivity against capacity is via synthetic tasks. The most common of these used to assess over-squashing are the RingTransfer tasks of Di Giovanni et al. (2023a). The goal of these tests is for a MPNN to ‘transfer’ features contained at a target node to a source node, across a large graph distance. Various graphs are tested, in particular a ring of nodes, but the common feature is that there exists a long shortest-path from the source to target node. All of these exhibit an exponentially growing receptive field of at least $2^{k}$ for $k$ layers, since each node is connected to at least two others; repeated aggregation and application of MPNN layers and nonlinearities makes this a good test of the sensitivity-based view of over-squashing.
However, this task is particularly ill-equipped to test the issue of storage capacity, as the only relevant information in the graph is that of the target node, and all intermediate nodes are assigned constant vectors of ones. In this way, there is only a single node’s representation worth of information to be transferred. It is unclear how much this task measures behavior found in real-world tasks: exponentially growing receptive fields will not be padded by nodes with identical representations. Figure 2 (left) visualizes the computational graph of RingTransfer, demonstrating that it is dominated by nodes containing no information. Therefore, this task exhibits a large computational bottleneck without any issues of saturating capacity: this highlights the fact that, beyond the computational bottleneck, saturated capacity is at least also dependent on the information content of the task.
Alon & Yahav (2021) introduced the Tree-NeighborsMatch task to measure capacity by propagating information from the leaf nodes of a variable-size binary tree. It shares similarities with the task we introduce in Section 3.2 in that it controls the amount of information that is forced into a single node representation. However, it propagates this information through a deep binary tree, requiring variable-depth MPNNs. This significantly harms sensitivity: we visualize Jacobian norms of a GCN acting on a deep binary tree vs a single layer tree with matching leaf counts in Figure 2, demonstrating that this sensitivity drops off far faster for deep GCNs. This is unsurprising given the bound of Equation 3: deep GCNs must additionally contend with “model” squashing terms of nonlinearity and weight contraction that scale exponentially with depth. Therefore performance degradation trends are due to both 1) saturating capacity and 2) low sensitivity; deep tasks such as Tree-NeighborsMatch are impacted by both over-squashing issues, rather than isolating the issue of capacity.
@@ 3.2 Neighbor Associative Recall: Isolating Storage Capacity
For a fixed neighborhood size $N$ we define a fixed set of keys and values, $N=|{\mathbb{K}}|=|{\mathbb{V}}|$, and a pair of learned vector embedding functions ${\bm{e}}_{k}:{\mathbb{K}}\to{\mathbb{R}}^{d_{\text{emb}}}$, ${\bm{e}}_{v}:{\mathbb{V}}\to{\mathbb{R}}^{d_{\text{emb}}}$ for embedding dimension $d_{\text{emb}}$. Each of the neighbor nodes $n$ has a different assigned key $k_{n}\in{\mathbb{K}}$, and also a value $v_{n}\in{\mathbb{V}}$, randomly sampled with replacement. The input feature vector of these nodes is a concatenation of the two learned embeddings ${\bm{x}}_{n}=[{\bm{e}}_{k}(k_{n});{\bm{e}}_{v}(v_{n})]\in{\mathbb{R}}^{2d_{\text{emb}}}$. The intermediate node and central node both have zero-valued feature vectors. Associated with the query node $q$ is a randomly sampled key-value node $m$; the input feature vector for the query node consists of the corresponding key embedding concatenated with a vector of zeroes, ${\bm{x}}_{q}=[{\bm{e}}_{k}(k_{m});\bm{0}]\in{\mathbb{R}}^{2d_{\text{emb}}}$. The model is trained such that the central node must predict the value $v_{m}$ associated with the sampled key node. Training is via cross-entropy loss where the target of the central node is a one-hot vector corresponding to a fixed ordering of $\mathbb{V}$}. This approach can be viewed as a graph adaptation of the sequence associative recall task of Schlag et al. (2021). Results are presented in Section 5.1. An alternative formulation of this task with a regression target is discussed in Section C.6.
@@
Prior work on over-squashing has focused almost exclusively on mitigating sensitivity issues, often through graph rewiring (Gasteiger et al., 2019; Gutteridge et al., 2023; Nguyen et al., 2023). Some work has implicitly tackled capacity over-squashing by moderating the flow of information into node representations (Bresson & Laurent, 2017; Finkelshtein et al., 2024; Errica et al., 2025) thus reducing capacity requirements, but we are unaware of any work that has attempted to increase capacity at an architecture level. Motivated by memory-capacity gains in sequence models (Ba et al., 2016; Beck et al., 2024), we introduce associative memory into an MPNN architecture to explicitly enlarge its information-storage capacity; we measure this in the graph setting using the NAR task introduced above. We further introduce the gating scheme of Beck et al. (2024) to investigate its efficacy in the graph setting, given strong sequence modeling performance. Since these adaptations are inspired in part by their successful use in xLSTM, we refer to our related graph architecture as gLSTM.
@@
The modified gLSTM update equations are given below. Highlighted in blue are the differences to xLSTM. Biases correspond exactly to xLSTM (Appendix B) and are omitted for clarity.
@@
Figure 4: gLSTM block structure. Gates shown in orange, query/key/value in dark blue. Aggr. represents aggregation across neighborhoods. Symbols $\odot,\otimes,+,\cdot$ denote Hadamard product, outer product, vector addition, matrix multiplication.
@@
This aggregation scheme appears to greatly improve gLSTM performance: our synthetic task in Section 5.1 significantly benefits from this aggregation scheme, and the ablations in Section C.2 demonstrate that it improves performance in all but one of the tested benchmarks. We hypothesize that – in addition to providing a highly connected computational graph that lessens over-squashing sensitivity bottleneck issues – this is because it also provides an extremely useful inductive bias for the recall mechanism of gLSTM. Information that has previously been stored in the associative memory is not then included in later message passing rounds, and later nodes are able to query this memory in isolation.
@@ 5.1 Neighbor Associative Recall
We train various models on NAR with varying neighbor count $N$, with results shown in Figure 5(a). Throughout this section we compare gLSTM using K-hop aggregation to GCN using standard aggregation, since gLSTM performs significantly better in this task when using K-hop aggregation whereas GCN performance is harmed by K-hop. We present additional results in Section C.5 where we separate by aggregation method and include results for a larger number of models. A comparison of the number of trainable parameters is shown in Figure 5(b). Fair comparison between matrix and vector memory is nontrivial, so we select these parameter counts to “favor” GCN.
@@
Figure 5: Test-set mean Accuracy (standard deviation shaded) for the NAR task, for gLSTM and GCN models with various hidden dimensions shown in Figure 5(a), number of trainable parameters in Figure 5(b). Note that gLSTM uses K-hop aggregation here, whereas GCN does not; see Section C.5 for separated performance by aggregation strategy.
@@ 5.2 How does Capacity Relate to Sensitivity?
We see therefore that sensitivity, as measured by the Jacobian norm, does not correlate with NAR performance. Given that NAR performance degradation is due to capacity over-squashing, we therefore observe that capacity over-squashing can occur without sensitivity over-squashing. This is clear from the fact that 1) sensitivity increases consistently for GCN models above $N=16$ to the point where it matches initial sensitivity, despite no increase in performance and 2) sensitivity for gLSTM tends to carry on increasing beyond where performance starts to degrade. We note these trends – as with all observations we make in this section – hold true for the NAR regression task in Section C.6.1.
@@ 5.3 Long Range Benchmarks
Table 1: Mean and standard deviation of $\log_{10}(\text{MSE})$, averaged over 4 random weight initializations on the GPP tasks from Gravina et al. (2023), from which we report baselines. See Section C.1 for discussion of baseline choice. Top score in bold, second underlined. Lower is better.
@@
gLSTM achieves comfortably state of the art results on the Diameter and Eccentricity GPP tasks, and very strong performance on SSSP; notably SSSP is the only tested task in which k-hop decreases performance. LRGB results show that gLSTM achieves strong performance in Peptides-Func but relatively weak performance on Peptides-Struct. We hypothesize that the weaker performance on Peptides-Struct may be due to long-range interactions being less relevant for this task, which is very effectively solved by a few-layer GCN. See Section C.2 for gLSTM ablations on these benchmarks and Section C.3 for details around hyperparameters used.